Shcrtly before midnight, on April 14, 1512, the Tranic collded with an iceberg and sank. Out of the 2190 passcngers, 1486 dicd thar nigh. The passcnos summarized data are shown below 320 285 TOS 880 2190 Round al answors to four decimals Supposc a passengcr is sclocicd at randcm: a) What is the probabity that the passcngor survivee? b) What is the prebability that the passenger was in 3rd class? What is the prababiity that the passcnger survvec? c) What is the probability that the passenger surived and was a crew mcmber? d) Wlhal is Ihe prctability at e passenger was rat in 1sl clas:s e) Gurviving is Click to sclectofbeing in 1st dass Suppisethe passenger seleclest was in 2d cass Whal is lhe pctetality Ihal the pasenge dient g) Wha is the probablity that a survhing passenger was a crew member? h) 1What is the prcbeatnity at the passenger sirvrver: or wsa re member Suring and being in 3rd clsareickto salc)
Answer #1

a)probability that passenger survived =705/2190=0.3219

b)P(3rd class)=705/2190=0.3219

c)P(survived and crew member)=207/2190=0.0945

d)P(not in 1st class)=1-320/2190=0.8539

e)not independent

f)P(died|2nd class)=164/285=0.5754

g)P(crew member|survived)=207/705=0.2936

h)P(survived or crew member)=(705+673)/2190=0.6292

i)not independent,